# Deep source separation of overlapping gravitational-wave signals and non-stationary noise artifacts

### v1

@@ I.1 Challenges in LISA data analysis
In contrast, MLE-based methods adopt a step-wise signal extraction process, estimating and subtracting sources like MBHBs, before analyzing GBs [23, 27]. This approach enables dynamic updates as new data arrives. However, uncertainty estimates are obtained from the Fisher information matrix, and thus are limited by the implicit assumption of Gaussian distributions without source correlations, though MCMC refinements can improve accuracy.
Given these challenges, research into complementary approaches for source separation in LISA data remains an active and evolving field. Deep-learning methods for data-driven feature extraction present a promising alternative by enabling direct source separation in a single step. These techniques offer advantages in computational efficiency, architectural flexibility, and scalability. The deep source separation approach holds potential for both low-latency applications and global-fit analyses.
@@ I.3 Contribution and overview of the paper
To address the challenge of source separation in LISA data, we introduce a deep learning-based framework designed to extract MBHBs, GBs, and instrumental glitches. Inspired by demucs (Deep extractor for music sources by Meta AI Research, see Ref. [37]), a model originally developed for the separation of musical instruments in audio data, our approach replaces the traditional sequential subtraction paradigm in LISA data analysis with single-step source extraction. By employing a shared encoder-decoder structure and latent space clustering, the model disentangles overlapping signals efficiently while remaining scalable to large source populations. A core feature of our framework is the frequency-binned output representation, which structures GB source separation on the basis of spectral content. This design helps mitigate source confusion by ensuring sources are dynamically clustered and disentangled in a learned feature space. Note that deep source separation operates independently of parameter estimation. This paper focuses on source separation, and parameter estimation based on its output is beyond the scope of this study.
@@ II.1 Massive black-hole binaries
As illustrated in Fig. 1, MBHBs with total masses between $10^{4}$ and $10^{7}$ $\mathrm{M}_{\odot}$ lie well within the LISA band, making them some of the loudest and most distant signals in LISA. Lower-mass binaries spend more time in the LISA band before merging, whereas higher-mass systems transition through LISA’s frequency range more quickly. Some MBHBs may also be multiband sources, appearing first in LISA before merging within the detection range of ground-based observatories.
@@ II.2 Galactic binaries
II.3 Extreme mass-ratio inspirals
EMRIs are another important class of sources expected in the LISA band, resulting from the inspiral of a compact object – typically a stellar-mass black hole, neutron star, or white dwarf – into a much more massive black hole, usually found at the center of a galaxy [45, 46, 47]. These systems generate long-lived, complex waveforms as the smaller object undergoes tens of thousands of orbits before merging. EMRIs encode precise information about the spacetime geometry of the central massive black hole, making them key probes for testing strong-field General Relativity and the nature of black holes. EMRIs emit gravitational waves in the 1 mHz to 10 mHz range, overlapping with the Galactic foreground and some lower-mass MBHB signals. Their waveforms are highly intricate, containing multiple harmonics that encode information about the mass, spin, and orbital eccentricity of the system. Unlike MBHBs, which evolve rapidly through the LISA band, EMRIs remain in LISA’s sensitivity window for months to years, requiring long-duration signal tracking and advanced matched-filtering techniques. While EMRIs are a key LISA science target, they are not included in the dataset considered for this paper. The same applies to stellar-origin black-hole binaries, and unmodeled gravitational wave bursts. Our focus is on MBHBs, GBs, and non-stationary noise artifacts. Future work needs to investigate incorporating these source types into deep source separation methods.
@@ II.4 Instrumental noise and glitches
One significant challenge is the presence of non-stationary glitches, short-duration noise transients caused by spacecraft systematics or environmental perturbations, such as micrometeoroid impacts. These glitches can mimic or obscure real gravitational-wave signals, making their identification and mitigation essential for accurate source separation [51, 52]. Characterizing instrumental noise is an active area of study, and techniques such as machine learning-based anomaly detection may play a crucial role in distinguishing true astrophysical signals from noise artifacts [53].
@@ II.5 Time-delay interferometry
Unlike ground-based detectors, which use simple Michelson interferometry, LISA’s evolving geometry introduces unique challenges in maintaining phase coherence, requiring an advanced signal-processing technique known as TDI [54, 55, 56]. TDI is designed to suppress laser frequency noise, which would otherwise overwhelm gravitational-wave signals. Laser noise suppression is accomplished by linearly combining and time-shifting LISA’s interferometric measurements to create virtual interferometers with equal arm lengths. It is important to note that the on-ground application of TDI transforms the data representation, altering the structure of signals compared to their raw strain measurements. For astrophysical data analysis, this transformation introduces additional complexity, since TDI modifies the apparent waveform morphology. In the context of machine learning, this requires that we train algorithms on TDI-processed data. Indeed, understanding these transformations is critical when designing traditional Bayesian inference or novel feature extraction methods.
@@ III Framework for deep source separation of LISA signals
The source separation framework presented in this paper is designed to extract astrophysical signals and to identify and estimate the evolution of glitches in the dataset. By taking advantage of the model’s ability to learn structured latent representations, we can distinguish glitches from genuine gravitational wave events, reducing the risk of misclassification. This capability is particularly valuable in scenarios where glitches overlap with astrophysical signals, ensuring that transient artifacts do not interfere with the accurate reconstruction of MBHBs or GBs.
@@ III.2.2 Latent representation and bottleneck transformation
Here, $I(A;B)$ represents the mutual information between two variables, quantifying how much knowing one reduces uncertainty about the other. The term $I(\tilde{z};x)$ ensures that the compressed representation $\tilde{z}$ still retains meaningful information about the input. The term $I(\tilde{z};n)$ represents the amount of irrelevant information (such as noise) that the bottleneck still encodes. The Lagrange multiplier $\beta$ controls the balance: a higher $\beta$ enforces stronger compression, forcing the network to discard more information. By optimizing Eq. 3 over the bottleneck parameters $\theta_{B}$, the model learns to remove redundancy while keeping only the most essential features, improving the separation of signals.
@@ III.2.3 Decoder
where $\hat{x}_{i}$ is the reconstructed output for the $i$-th source, and $D_{i}(\,\cdot\,;\theta_{D_{i}})$ represents the decoder network parameterized by $\theta_{D}$, responsible for reconstructing individual components. The decoder can apply a series of transposed convolutions to progressively upsample and restore temporal structures from the compressed latent representation. To enhance reconstruction accuracy, skip connections can be incorporated, allowing the network to retain fine-grained details by reintroducing relevant features from earlier encoding layers. Typically, source-specific activation functions are employed to ensure that each decoder head reconstructs only its assigned target, preventing interference between different signal types.
@@ III.3 Demucs as an example of an established encoder-decoder model
The encoder in demucs comprises multiple convolutional layers that progressively downsample the input, capturing hierarchical features. Each convolutional block integrates standard convolutions, batch normalization, and nonlinear activation functions to enhance feature extraction. To further improve its capacity to model long-range temporal dependencies, demucs incorporates BLSTM layers within the bottleneck, allowing the model to process both forward and backward temporal information. This feature is particularly advantageous for disentangling overlapping musical components, which frequently exhibit strong temporal correlations. Additionally, demucs employs gated linear units as activation functions, which enhance the model’s expressiveness by selectively regulating information flow.
@@ III.4 Modifying demucs for application in LISA
Adapting demucs for LISA data requires modifications that account for the unique characteristics of gravitational-wave signals. Unlike conventional audio streams, LISA data comprises a superposition of overlapping astrophysical waveforms and transient instrumental glitches. The primary challenge in this adaptation lies in accurately isolating individual sources while ensuring the scalability of the separation model. To address this, we design a modified architecture to structure the separation task across three dedicated decoder heads following a shared encoder and bottleneck representation. The first decoder is responsible for reconstructing MBHB signals, assuming for simplicity that individual MBHB events do not overlap in time. The second decoder isolates non-stationary noise artifacts, such as transient glitches. The third decoder is specifically designed for GBs and consists of multiple output channels, each corresponding to a predefined frequency bin. This multi-output design enables the model to disentangle individual GB sources while maintaining scalability. Instead of assigning a separate decoder to each GB source – an approach that quickly becomes computationally prohibitive – the frequency space is discretized into small bins, ensuring that each bin contains at most one dominant source. In our framework, we intentionally omit skip connections to maintain simplicity. As a result, our bottleneck applies a weaker compression compared to demucs. Details will follow.
We use LISA’s TDI data streams A, E, and T [72] as input to the model. For now, each TDI channel is processed independently, applying the same separation model. An extended version of the framework could allow multiple TDI streams to be processed in parallel. This extension might further optimize source separation accuracy by taking advantage of complementary information from different TDI observables during training and inference, but it is not required for this demonstration. The schematic architecture of the demucs-inspired multi-source extraction model is depicted in Fig. 2.
@@ III.4.3 Loss calculation and training
$L_{\text{Total}}=L_{\text{MBHB}}+L_{\text{GB}}+L_{\text{glitch}},$ (5)
where $L_{\text{MBHB}}$, $L_{\text{GB}}$, and $L_{\text{glitch}}$ represent the reconstruction losses for each source type.
@@
For GB sources, a frequency-binned representation is used to ensure structured separation of the overlapping signals in the spectral domain. We map the GB signals into discrete frequency bins using a bin assignment function, which finds the nearest predefined frequency bin for each GB signal in the data and accumulates its contribution into the corresponding bin. Formally, for a batch of GB signals $x_{\text{GB}}$ with associated frequencies $f_{\text{GB}}$, the binned representation is given by:
$&x_{\text{GB}}^{\text{bin}}[b,b_{i}]=\sum_{i}x_{\text{GB}}[b,i]\\ &\text{with}\quad b_{i}=\arg\min|f_{\text{GB},i}-f_{\text{bins}}|\end{aligned}$
where $b$ represents the batch index, $b_{i}$ is the index of the closest frequency bin for source $i$, and $f_{\text{bins}}$ is a predefined list of frequency bin edges used to categorize GB signals based on their frequencies. The frequency-binned ground truth signal is then compared to the predicted GB output using MSE:
@@
Given that LISA observations span months to years, the model must handle long-duration signals while maintaining source coherence. Training on full-time-series data is computationally infeasible, so the model is trained on short-duration snippets analogously to the approach followed in demucs. During inference, overlapping segments are processed independently, and efforts are currently underway to develop and refine a method for seamlessly merging these segments using a weighted averaging scheme. This work will be presented in a follow-up paper.

### v2

@@ I.1 Challenges in LISA data analysis
In contrast, MLE-based methods typically follow a deterministic, step-wise signal extraction strategy, where sources such as MBHBs are estimated and subtracted before proceeding to fainter components like GBs [23, 27]. While hierarchical subtraction is also employed in MCMC-based pipelines to enhance sampling efficiency, it is integrated within a broader Bayesian framework that jointly estimates all sources and parameters.
Given these challenges, research into complementary approaches for source separation in LISA data remains an active and evolving field. Deep-learning methods for data-driven feature extraction present a promising alternative by enabling direct source separation in a single step. These techniques offer advantages in computational efficiency, architectural flexibility, and scalability. Related work in the context of ground-based detectors, such as DeepExtractor [28], has demonstrated the potential of deep learning for reconstructing gravitational-wave signals and mitigating transient noise artifacts. Moreover, UnMixFormer [29] has demonstrated the effectiveness of attention-based architectures for counting and separating overlapping compact binary coalescence signals in ground-based detector data. Besides, simulation-based inference methods, such as Sequential Neural Likelihood [30], have recently been applied to LISA MBHB signals, enabling efficient posterior estimation with fewer simulator calls than traditional MCMC.
A key motivation for dedicated source separation and reconstruction stems from the fact that many gravitational-wave signals in LISA data overlap in both time and frequency, leading to strongly blended mixtures in the recorded data streams. This overlap poses a major obstacle for traditional Bayesian inference: the resulting likelihood surface becomes highly multimodal and degenerate, especially when multiple signals occupy the same frequency band. For example, accurately characterizing a faint GB becomes significantly more difficult when its signal is masked by a nearby, louder source – whether of the same class or a different type. Without some form of source separation, classical parameter estimation methods must attempt to jointly fit overlapping signals, a process that is computationally expensive and scales poorly with source density.
Deep source separation addresses this problem by disentangling overlapping signals before parameter inference. This approach can transform the inference pipeline from a monolithic global fit into a modular two-stage process: (1) extract individual sources from the mixture and (2) perform parameter estimation on each extracted source independently or in smaller batches. As a result, source separation simplifies the inference landscape, reduces the dimensionality of the search space, and enables scalable parallelization.
@@ I.3 Contribution and overview of the paper
To address the challenge of source separation in LISA data, we introduce a deep learning-based framework designed to extract MBHBs, GBs, and instrumental glitches. Inspired by demucs (Deep extractor for music sources by Meta AI Research, see Ref. [40]), a model originally developed for the separation of musical instruments in audio data, our approach replaces the traditional sequential subtraction paradigm in LISA data analysis with single-step source extraction. While we do not reuse any code from demucs, we adopt a very similar architectural design. By employing a shared encoder-decoder structure and latent space clustering, the model disentangles overlapping signals efficiently while remaining scalable to large source populations. A core feature of our framework is the frequency-binned output representation, which structures GB source separation on the basis of spectral content. This design helps mitigate source confusion by ensuring sources are dynamically clustered and disentangled in a learned feature space. Note that deep source separation operates independently of parameter estimation. This paper focuses on source separation, and parameter estimation based on its output is beyond the scope of this study.
@@ II.1 Massive black-hole binaries
As illustrated in Fig. 1, MBHBs with total masses between $10^{4}$ and $10^{7}$ $\mathrm{M}_{\odot}$ lie well within the LISA band, making them some of the loudest and most distant signals in LISA. Lower-mass massive black hole binaries (MBHBs) emit gravitational waves at higher frequencies and thus spend more time evolving within the LISA band before merger. In contrast, higher-mass systems have a lower merger frequency, often exiting the LISA sensitivity range before reaching its upper end, resulting in shorter in-band durations. Some MBHBs may also be multiband sources, entering the LISA band years before merger and later merging within the sensitivity window of ground-based detectors
@@ II.2 Galactic binaries
While this foreground is often modeled as stationary over short durations, it is in fact non-stationary on mission timescales. Two main mechanisms introduce this temporal evolution: (i) the intrinsic frequency drift of individual binaries due to gravitational radiation reaction, and (ii) the periodic Doppler modulation induced by LISA’s orbital motion around the Sun. These effects cause the apparent frequency and amplitude of sources to vary over time, imprinting slowly changing patterns on the composite foreground signal. On long timescales, such modulations can help distinguish overlapping sources by introducing characteristic time-frequency signatures that aid in source identification. Capturing these non-stationary features in data-driven models requires a large and diverse training sets that reflects the full range of time-dependent behavior expected during the mission.
II.3 Sources beyond the present study
EMRIs are another important class of sources expected in the LISA band, resulting from the inspiral of a compact object – typically a stellar-mass black hole, neutron star, or white dwarf – into a much more massive black hole, usually found at the center of a galaxy [48, 49, 50]. These systems generate long-lived, complex waveforms as the smaller object undergoes tens of thousands of orbits before merging. EMRIs encode precise information about the spacetime geometry of the central massive black hole, making them key probes for testing strong-field General Relativity and the nature of black holes. EMRIs emit gravitational waves in the 1 mHz to 10 mHz range, overlapping with the Galactic foreground and some lower-mass MBHB signals. Their waveforms are highly intricate, containing multiple harmonics that encode information about the mass, spin, and orbital eccentricity of the system. Unlike MBHBs, which evolve rapidly through the LISA band, EMRIs remain in LISA’s sensitivity window for months to years. Accurately detecting, separating, and reconstructing such signals may require several methodological extensions to the framework presented in this paper, including, for example, hierarchical or multi-resolution network architectures, memory-aware encoders, or recurrent modules capable of capturing long-term temporal dependencies. Switching from raw time-domain inputs to time-frequency representations may also be beneficial, as they offer a more compact and structured view of slowly evolving signals like EMRIs. Additionally, due to their typically low signal-to-noise ratios, EMRIs are expected to require substantially larger training datasets to achieve reliable separation and reconstruction, which stands in contrast to the limited-data, proof-of-concept setting considered in this work. For these reasons, shared with current global-fit analyses that similarly omit EMRIs, we do not include them in the present study. The same applies to stellar-origin black hole binaries and unmodeled gravitational-wave bursts. Our current focus is on MBHBs, GBs, and non-stationary noise artifacts. Exploring the necessary architectural and data-driven adaptations remains an important direction for future research and will be essential to extending deep source separation methods to these challenging classes of sources.
@@ II.4 Instrumental noise and glitches
One significant challenge is the presence of glitches, short-duration noise transients caused by spacecraft systematics, or environmental perturbations, such as micrometeoroid impacts. These glitches can mimic or obscure real gravitational-wave signals, making their identification and mitigation essential for accurate source separation [54, 55]. Characterizing instrumental noise is an active area of study, and techniques such as machine learning-based anomaly detection may play a crucial role in distinguishing true astrophysical signals from noise artifacts [56].
@@ II.5 Time-delay interferometry
Unlike ground-based detectors, which use simple Michelson interferometry, LISA’s evolving geometry introduces unique challenges in maintaining phase coherence, requiring an advanced signal-processing technique known as TDI [57, 58, 59]. TDI is designed to suppress laser frequency noise, which would otherwise overwhelm gravitational-wave signals. Laser noise suppression is accomplished by linearly combining and time-shifting LISA’s interferometric measurements to create virtual interferometers with equal arm lengths. It is important to note that the on-ground application of TDI transforms the data representation, altering the structure of signals compared to their raw strain measurements. In the context of machine learning, this requires that we train algorithms on TDI-processed data. Indeed, understanding these transformations is critical when designing traditional Bayesian inference or novel feature extraction methods.
@@ III Framework for deep source separation of LISA signals
The source separation framework presented in this paper is designed to extract astrophysical signals and to identify and estimate glitches in the dataset. By taking advantage of the model’s ability to learn structured latent representations, we can distinguish glitches from genuine gravitational wave events, reducing the risk of misclassification. This capability is particularly valuable in scenarios where glitches overlap with astrophysical signals, ensuring that transient artifacts do not interfere with the accurate reconstruction of MBHBs or GBs.
@@ III.2.2 Latent representation and bottleneck transformation
Here, $I(A;B)$ denotes the mutual information between variables $A$ and $B$, quantifying how much knowing one reduces uncertainty about the other. The term $I(\tilde{z};x)$ ensures that the compressed representation retains meaningful information about the input, while $I(\tilde{z};n)$ penalizes the retention of irrelevant information. In our context, $n$ corresponds to components of the input that are not meant to be explicitly reconstructed, primarily the quasi-stationary instrumental noise. To guide the encoder and bottleneck toward discarding irrelevant components, we use a frequency-dependent noise model to generate diverse time-domain noise realizations during training. This exposure enables the network to distinguish between signals of interest and stationary noise, and to focus its representational capacity on features relevant to signal reconstruction.
It remains an open question how performance is affected when the evaluation data exhibits noise properties that differ from the training distribution. Note that this is not a specific limitation of our method, but a general challenge in machine learning-based analyses of noisy, high-dimensional measurements.
@@ III.2.3 Decoder
where $\hat{x}_{i}$ is the reconstructed output for the $i$-th source, and $D_{i}(\,\cdot\,;\theta_{D_{i}})$ represents the decoder network parameterized by $\theta_{D_{i}}$, responsible for reconstructing individual components. The decoder can apply a series of transposed convolutions to progressively upsample and restore temporal structures from the compressed latent representation. To enhance reconstruction accuracy, skip connections can be incorporated, allowing the network to retain fine-grained details by reintroducing relevant features from earlier encoding layers. Typically, source-specific activation functions are employed to ensure that each decoder head reconstructs only its assigned target, preventing interference between different signal types.
@@ III.3 Demucs as an example of an established encoder-decoder model
The encoder in demucs comprises multiple convolutional layers that progressively downsample the input, capturing hierarchical features. Each convolutional block integrates standard convolutions, batch normalization, and nonlinear activation functions to enhance feature extraction. To improve its capacity to model temporal dependencies within each input segment, demucs incorporates BLSTM layers within the bottleneck. These allow the model to process both forward and backward temporal context over the segment duration. This is particularly advantageous for disentangling overlapping musical components that exhibit strong temporal structure – such as harmonically related instruments or time-aligned effects – within the scope of each training snippet. Additionally, demucs employs gated linear units as activation functions, which enhance the model’s expressiveness by selectively regulating information flow.
@@ III.4 Modifying demucs for application in LISA
Adapting demucs for LISA data requires modifications that account for the unique characteristics of gravitational-wave signals. Unlike conventional audio streams, LISA data comprises a superposition of overlapping astrophysical waveforms and transient instrumental glitches. The primary challenge in this adaptation lies in accurately isolating individual sources while ensuring the scalability of the separation model. To address this, we design a modified architecture to structure the separation task across three dedicated decoder heads following a shared encoder and bottleneck representation. The first decoder is responsible for reconstructing MBHB signals, assuming for simplicity that individual MBHB events do not overlap in time. The second decoder isolates non-stationary noise artifacts, such as transient glitches. The third decoder is specifically designed for GBs and consists of multiple output channels, each corresponding to a predefined frequency bin. This multi-output design enables the model to disentangle individual GB sources while maintaining scalability. Instead of assigning a separate decoder to each GB source – an approach that quickly becomes computationally prohibitive – the frequency space is discretized into small bins, ensuring that each bin contains at most one dominant source. This mirrors the assumption used for MBHBs in the time domain, where each time chunk is assumed to contain at most one MBHB signal. In our framework, we intentionally omit skip connections to maintain simplicity. As a result, our bottleneck applies a weaker compression compared to demucs. Details will follow.
We use LISA’s TDI data streams A, E, and T [75] as input to the model. These channels form an orthogonal and widely used basis that spans the full gravitational-wave response space of the detector; any other complete set of TDI combinations would be equally valid from an information perspective. In the current setup, each TDI channel is processed independently using a shared separation model. This simplification reduces model complexity and training cost, but it discards potentially useful cross-channel correlations. As a result, source reconstructions across the channels may become inconsistent in the presence of non-stationary noise or low signal-to-noise ratios. Future extensions of the framework will adopt joint multi-channel processing – such as through shared or cross-channel encoder structures – to better leverage the complementarity of different TDI observables. While such modifications may improve reconstruction accuracy and consistency, they are not essential for the proof-of-concept separation task presented here. Ultimately, training the network to recover the underlying gravitational-wave strain in the barycentric frame, e.g., the $h_{+}$ and $h_{\times}$ polarizations, could offer further advantages for downstream parameter estimation and is left for future work. The schematic architecture of the demucs-inspired multi-source extraction model is depicted in Fig. 2.
@@ III.4.3 Loss calculation and training
$L_{\text{Total}}=L_{\text{MBHB}}+L_{\text{glitch}}+L_{\text{GB}},$ (5)
where $L_{\text{MBHB}}$, $L_{\text{glitch}}$, and $L_{\text{GB}}$ represent the reconstruction losses for each source type.
@@
For GB sources, a frequency-binned representation is used to ensure structured separation of the overlapping signals in the spectral domain. Each GB signal is assigned to its nearest frequency bin based on its central frequency. Formally, for each batch $b$ and bin index $k$, the binned GB signal is computed as
$x_{\text{GB}}^{\text{bin}}[b,k,t]=\sum_{i\in\mathcal{I}_{b,k}}x_{\text{GB}}[b,i,t],$ (8)
where $x_{\text{GB}}[b,i,t]$ denotes the waveform of the $i$-th GB source and
$\mathcal{I}_{b,k}=\left\{i\,\middle|\,\arg\min_{j}\left|f_{\text{GB}}[b,i]-f_{\text{bins}}[j]\right|=k\right\}$ (9)
is the set of sources assigned to bin $k$ based on their frequency proximity to the predefined bin centers $f_{\text{bins}}$. The frequency-binned ground truth signal is then compared to the predicted GB output using MSE:
@@
Given that LISA observations span months to years, the model must be capable of processing long-duration signals while preserving temporal coherence. However, training directly on full-length time-series data is computationally prohibitive due to the memory and processing demands of handling inputs with millions of time steps. Such long sequences would exceed the memory limits of standard hardware, particularly when used in conjunction with deep convolutional encoder-decoder architectures.
To address this, we adopt an approach similar to that used in demucs, where the model is trained on randomly sampled short-duration segments (e.g., minutes to hours). This strategy significantly reduces memory usage, allows for efficient batch processing, and promotes generalization. When applying the model to full-length LISA data, it will be necessary to divide the time series into overlapping segments, process each independently, and then merge the outputs using a weighted averaging scheme. This stitching mechanism – also employed in demucs – ensures continuity across segment boundaries and mitigates edge artifacts. This work will be presented in a follow-up paper.
@@
III.5 Future directions for resolving overlapping Galactic binaries
The current GB decoder prototype is based on a simplifying assumption: that the frequency axis can be discretized into narrow bins such that each bin contains at most one GB signal. In practice, the spectral density of the GB population will be high, especially at low frequencies, making this assumption increasingly fragile. In densely populated regions of the spectrum, multiple GB sources with closely spaced frequencies may fall within the same bin, even under carefully optimized decoder binning schemes. These near-degenerate signals, though individually narrow-band, can interfere constructively or destructively, particularly when their amplitudes are similar, posing a fundamental challenge for source separation in LISA. While differences in parameters such as sky position can, in principle, induce distinct Doppler and amplitude modulation patterns that aid disentanglement, the current purely bin-wise decoder design lacks the capacity to fully exploit such subtle variations. As a result, the model may struggle to accurately resolve individual signals in frequency-overlapping scenarios – even if their SNRs would allow distinguishability in theory.
A full treatment of this issue lies beyond the scope of the present study. Nevertheless, we outline some architectural directions that move beyond the current one-source-per-bin assumption. One promising strategy is to adopt a dynamic multi-slot decoding scheme, in which each frequency bin produces a flexible number of output slots, with each slot representing a distinct candidate source. Rather than fixing the number of slots a priori, mechanisms such as Slot Attention [77] or set-based transformers could iteratively infer both the number and identity of sources present, conditioned on local and global features. This approach would allow the model to adaptively allocate representational capacity based on local source density. To better resolve sources whose features span multiple bins (e.g., due to modulation or frequency drift), future architectures could also incorporate sequence modeling across bins, using temporal convolution, recurrent layers, or attention-based modules. These models can exploit correlations between neighboring bins to disentangle overlapping signals that cannot be separated using purely local information. Training such models would require a combination of reconstruction loss, permutation-invariant supervision (e.g., using the Hungarian algorithm [78]), and disentanglement-promoting regularizers (such as orthogonality penalties or contrastive objectives).
Future work will evaluate the effectiveness of this strategy on realistically dense GB populations, where overlapping sources are not rare exceptions but rather a fundamental aspect of the data.